Car rental agency A will rent a compact car for $35 per day and an additional charge of $.24 per mile. Car rental agency B will charge only $.16 per mile but charges $41 per day. If Adam wanted to rent a car for three days, how many miles would Adam have to drive to make car rental agancy B a beter bargain?
Answer by checkley77(12569)

3*35+.24x=3*41+.16x
105+.24x=123+.16x
.24x-16x=123-105
.08x=18
x=18/.08
x=225 miles driven would make the 2 rentals the same for the 3 days.
Proof:
105+.24*225=123+.16*225
105+54=123+36
159=159